OpenStudy (naveenbhatia1312):

A person is watching a boat from the top of a lighthouse. The boat is approaching the lighthouse directly. When first noticed the angle of depression to the boat is 16° 23'. When the boat stops, the angle of depression is 49° 29'. The lighthouse is 200 feet tall. How far did the boat travel from when it was first noticed until it stopped? Round to the nearest hundredth. (5 points) 554.56 ft 531.86 ft 494.06 ft 509.36ft

OpenStudy (anonymous):

This is not difficult. Use the tangent function of each angle. tan Θ = height of lighthouse / distance of boat from lighthouse. tan (15.85°) = 200 / x x = 200 / tan 15.85° x ≈ 704.438 ft tan (52.2°) = 200 / x' x' = 200 / tan (52.2°) x' ≈ 155.136 ft.. The distance the boat traveled is equal to x - x': x - x' = 704.438 ft - 155.136 ft = 549.302 ft. We can then round the answer to 549.30 ft.

OpenStudy (anonymous):



OpenStudy (anonymous):

did that help?

OpenStudy (naveenbhatia1312):

I got 549.30 as well but im confused because its not one of the answer choices

OpenStudy (anonymous):

what are the answer choices?

OpenStudy (naveenbhatia1312):

554.56 ft 531.86 ft 494.06 ft 509.36ft

OpenStudy (anonymous):

oh ummmm let me see with @zepdrix

OpenStudy (naveenbhatia1312):

ok cool

zepdrix (zepdrix):

degrees and minutes? 0_o ooo this one is annoying.. thinking

zepdrix (zepdrix):

hmm i got one of the answer choices :3 boy this one is tough to explain though...

zepdrix (zepdrix):